Solve each system by the addition method.
x+ y=13
22x+36y=356

Answer by checkley71(8403)

MULTIPLY THE FIRST EQUATION BY -22 & ADD:
-22X-22Y=-286
22X+36Y=356
--------------------------
14Y=70
Y=70/14
Y=5 ANSWER.
X+5=13
X=13-5
X=8 ANSWER.
PROOF
22*8+36*5=356
176+180=356
356=356